# Sparse-grid polynomial interpolation approximation and integration for parametric and stochastic elliptic PDEs with lognormal inputs

### v14

@@ 6 Elliptic PDEs with affine inputs
Let Assumption III hold. Let $a=b$ for the Jacobi probability measure $\nu_{a,b}({\boldsymbol{y}})$, and the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{\rm ev}(\xi)$ be the set defined by (3.10) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then for the quadrature operator ${\mathcal{Q}}_{G_{\rm ev}(\xi)}$ generated by the interpolation operator ${\mathcal{I}}^{*}_{G_{\rm ev}(\xi)}:{\mathcal{L}}_{2}(W^{r})\to{\mathcal{V}}(G_{\rm ev}(\xi))$, we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}u\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.6)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.7)
@@
Let $a=b$ for the Jacobi probability measure $\nu_{a,b}({\boldsymbol{y}})$, and the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $\Lambda_{\rm ev}(\xi)$ be the set defined by (3.45) for $\sigma_{\boldsymbol{s}}:=\beta_{1;{\boldsymbol{s}}}$ as in (6.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (6.8)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (6.9)

### v15

@@ 6 Elliptic PDEs with affine inputs
Let Assumption III hold. Let $a=b$ for the Jacobi probability measure $\nu_{a,b}({\boldsymbol{y}})$, and the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{\rm ev}^{*}(\xi)$ be the set defined by (5.6) for $q_{2}=q_{r}$ and for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then for the quadrature operator ${\mathcal{Q}}_{G_{\rm ev}^{*}(\xi)}$ generated by the interpolation operator ${\mathcal{I}}^{*}_{G_{\rm ev}^{*}(\xi)}:{\mathcal{L}}_{2}(W^{r})\to{\mathcal{V}}(G_{\rm ev}^{*}(\xi))$, we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}^{*}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}u\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{r}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{r}-1/2.\end{cases}$ (6.6)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}^{*}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}^{*}(\xi_{n})}\langle\phi,u\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 2/q_{r}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>2/q_{r}-1/2.\end{cases}$ (6.7)
@@
Let $a=b$ for the Jacobi probability measure $\nu_{a,b}({\boldsymbol{y}})$, and the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ with some $0<q_{1}=q<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $\Lambda_{\rm ev}^{*}(\xi)$ be the set defined by (5.10) for $\sigma_{\boldsymbol{s}}:=\beta_{1;{\boldsymbol{s}}}$ as in (6.1). Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}^{*}(\xi_{n}))|\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}^{*}(\xi_{n})}u\right\|_{V}\ \leq\ Cn^{-(2/q-1/2)}.$ (6.8)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}^{*}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}^{*}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (6.9)